OpenStudy (anonymous):

Subtract and give the answer in simplest form 3x/x-5 - 5/5-x

OpenStudy (anonymous):

You start with\[\frac{3x}{x-5}-\frac{5}{5-x}\] Notice that you can rewrite \[-\frac{5}{5-x}\]as\[\frac{5}{x-5}\] So you have\[\frac{3x}{x-5}+\frac{5}{x-5}\]

OpenStudy (anonymous):

\[3x/(x-5) - 5/(5-x) = 3x/(x-5) + 5/(x-5) = (3x + 5)/(x-5)\]